OpenStudy (jimmiejones153):

4x + 1 over 3y2 What is the value of the expression above when x = 2 and y = 3? You must show all work and calculations to receive full credit. i need to get this done before 9:50

sam (.sam.):

Do you know substitution?

OpenStudy (jimmiejones153):

yes

sam (.sam.):

So \[4x+\frac{1}{3y^2}\] When x=2, y=3, \[4(2)+\frac{1}{3(3)^2}\] What do you get?

OpenStudy (jimmiejones153):

idk

sam (.sam.):

What is 4 times 2?

OpenStudy (jimmiejones153):

8

sam (.sam.):

Now for the denominator, the square of 3?

OpenStudy (jimmiejones153):

what

sam (.sam.):